OpenStudy (anonymous):

Listed below are brain volumes (cm3) of unrelated subjects used in a study. Assume the data is from a simple random sample. Use a 0.01 significance level and the critical values method to test the claim that the population of brain volumes has a mean equal to 1100.0cm3. Be sure to include all steps of a hypothesis test. 963 1027 1272 1079 1070 1173 1067 1347 1100 1204

OpenStudy (amistre64):

what steps have you got so far?

OpenStudy (anonymous):

I have the critical values of z= -2.575 and z= 2.575

OpenStudy (anonymous):

And the formula z= xbar-u/sigma/square root of n

OpenStudy (anonymous):

I am not sure if that is the right formula to use

OpenStudy (amistre64):

that formulas fine hmm, all thats left it appears is to find the mean of the sample, and possibly the number of dataelements in the sample.

OpenStudy (anonymous):

I am not sure how to find sample mean and standard deviation

OpenStudy (amistre64):

not to sure how your material defines 'all steps' but in general, we state a hypothesis. then compare the results to determine which hypot is most likely you seem to have the right idea that its a 2 tailed test of equality. Ho: u = 1100 Ha: u not= 1100 then develop the CI such that:\[\pm2.575=\frac{\bar x-\mu}{\sigma/\sqrt n}\] \[CI:~\bar x\pm2.575({\sigma/\sqrt n})=\mu\]

OpenStudy (amistre64):

standard deviation is given in the information .... and if you dont know how to find the mean of a set of numbers by now, then you might need to review the first chapter of your material

OpenStudy (amistre64):

a little error in my thought, i was confusing a precious confidence interval into this setup :)

OpenStudy (anonymous):

Okay I know how to find the mean it is just adding the numbers up and dividing but i wasnt sure if thats what it meant by " sample mean"

OpenStudy (amistre64):

the data set provided is a sample ... so the sample mean is the mean of the sample data.

OpenStudy (anonymous):

Okay thank you, but I still am not sure what the standard deviation is?

OpenStudy (amistre64):

what does the information tell us it is?

OpenStudy (amistre64):

youre right ... im on a different problem in my head still

OpenStudy (anonymous):

I don't know... It tells us the significance level and tells us the population mean

OpenStudy (amistre64):

since there is not population sd given, we can subsitute in the sample sd, and use a t distribution. does that sound familiar?

OpenStudy (anonymous):

sd?

OpenStudy (amistre64):

\[s=\frac{\sum(x-\bar x)^2}{n-1}\]

OpenStudy (amistre64):

standard deviation (sd)

OpenStudy (anonymous):

is that an 8?

OpenStudy (amistre64):

...no sigma is the variable used for a population sd s is the variable used to a sample sd just like: mu is used for a population mean, and xbar is used for a sample mean

OpenStudy (amistre64):

thats an s

OpenStudy (anonymous):

Thankyou

OpenStudy (amistre64):

if sigma (population sd) is known\[\large z=\frac{\bar x-\mu}{\sigma/\sqrt n}\] if sigma is NOT known, we can substitute the sample sd, s, which gives out setup a t distribution instead of a normal distribution: \[\large t=\frac{\bar x-\mu}{s/\sqrt n}\]

OpenStudy (amistre64):

using excel, we can quickly find the sample mean, variance, and sd

OpenStudy (anonymous):

Woah I didn't know you could do that, Haha!

OpenStudy (amistre64):

using a t chart .... with a 2 tailed significant level of .01, and dF of 10-1 = 9 http://web.pdx.edu/~stipakb/download/PA551/t-distribution_table.gif t = 3.25

OpenStudy (amistre64):

if out test statisitic of:\[\frac{1130.2-1100.0}{117.45/\sqrt{10}}>3.25\]then we fall in the tails of our hypothesis setup. and reject the null hypot.

OpenStudy (anonymous):

How did you get 10-1?

OpenStudy (amistre64):

there are 10 data elements listed. so for a t distribution we would need to determine the degrees of freedom as n-1

OpenStudy (amistre64):

have you covered t distributions?

OpenStudy (anonymous):

No

OpenStudy (amistre64):

hmm, yet this is what the information would call for. its similar to a normal distribution, and as the number of data elements in the sample increases it better approximates the normal distribution. its used in the same manner but with a few tweaks

OpenStudy (anonymous):

I am taking an online statistics class, and the modules we have only cover certain things and we are left with the textbook. I am struggling to teach myself because there are very few examples in the textbook

OpenStudy (amistre64):

ive used this site to learn a lot of stuff and to teach myself as well. just graduated with a bachelors in math :)

OpenStudy (anonymous):

Good for you!

OpenStudy (amistre64):

yeah ... now i just need a job lol http://stattrek.com/probability-distributions/t-distribution.aspx this sites readable to me and covers the t distribution rather simply.

OpenStudy (anonymous):

So it says to not use the t-distribution with sample sizes less than 40?

OpenStudy (amistre64):

i read that ..... im gonna do another search.

OpenStudy (amistre64):

heres a more simplified rule lol

OpenStudy (amistre64):

that doc was from web.pdx.edu